Please help me with this, I have no idea how to do find out the moles of acid/moles of glucose fermented.

Source

Gas in Durham’s Tube

Benedict’s Test (Color)

Benedict’s Test Interpretation

Titration (mL)titrant

Moles of acid/moles of glucose fermented

Control

-

orange

1-1.5% reducing sugar

13.2mL

Homofermentative Lactobacillus

-

green

0.1-0.5% reducing sugar

75.9mL

Heterofermentative

Lactobacillus

+

green

0.1-0.5% reducing sugar

129.1mL

Subtract the volume of 0.1 N NaOH titrant used to titrate the uninoculated control from the volume of 0.1 N NaOH titrant used to titrate each of the inoculated cultures. Assuming that the remaining titratable acid is lactic acid produced via fermentation, calculate the moles of acid produced per mole of glucose fermented and compare the values determined for each culture.
Answer #1

Concept: Moles of acid/moles of glucose fermented = the moles of the titrant (NaOH).

For NaOH solution, normality = molarity because equivalent weight = molecular weight

(i) Moles of acid/moles of glucose fermented = (13.2/1000) L * 0.1 mol/L = 0.00132 mol

(ii) Moles of acid/moles of glucose fermented = (75.9/1000) L * 0.1 mol/L = 0.00759 mol

(ii) Moles of acid/moles of glucose fermented = (129.1/1000) L * 0.1 mol/L = 0.01291 mol